# YOLO11 to Its Genesis: A Decadal and Comprehensive Review of The You Only Look Once (YOLO) Series

### v5

@@ Abstract
Given the rapid emergence and applications of Large Language This review systematically examines the progression of the You Only Look Once (YOLO) object detection algorithms from YOLOv1 to the recently unveiled YOLO11 (or YOLOv11). Employing a reverse chronological analysis, this study examines the advancements introduced by YOLO algorithms, beginning with YOLOv11 and progressing through YOLOv10, YOLOv9, YOLOv8, and subsequent versions to explore each version’s contributions to enhancing speed, detection accuracy, and computational efficiency in real-time object detection. By detailing the incremental technological advancements in subsequent YOLO versions, this review chronicles the evolution of YOLO, and discusses the challenges and limitations in each earlier versions. The evolution signifies a path towards integrating YOLO with multimodal, context-aware, and Artificial General Intelligence (AGI) systems for the next YOLO decade, promising significant implications for future developments in AI-driven applications.
@@
You Only Look Once, YOLO, YOLOv1 to YOLOv11, YOLO configurations, CNN, Deep learning, Real-time object detection, Artificial intelligence, Computer vision, Healthcare and Medical Imaging, Autonomous vehicles, Traffic safety, Industrial manufacturing, Surveillance, Agriculture
@@ I-D You Only Look Once Approach
The ”You Only Look Once” (YOLO) object detection algorithm was first introduced by Joseph Redmon et al.,[58] in 2016, revolutionized real-time object detection by combining region proposal and classification into a single neural network, significantly reducing computation time. YOLO’s unified architecture divides the image into a grid, predicting bounding boxes and class probabilities directly for each cell, enabling end-to-end learning[58]. YOLO is versatile, and its real-time detection capabilities have revolutionized autonomous vehicles and traffic safety[4, 59, 16, 60], healthcare [61, 62, 63], industrial applications[64, 65, 59], surveillance and security[66, 67] and agriculture [68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78], where accuracy and speed are crucial.
In autonomous vehicles and traffic safety, YOLO has been widely adopted to improve safety in both road transport and aviation, allowing real-time object detection for collision avoidance, traffic monitoring, pedestrian detection, traffic sign detection, and aviation hazard analysis. This application aids in minimizing accidents and improving operational efficiency[4, 79, 59].
In healthcare, YOLO has been instrumental in assisting and improving diagnostic processes and treatment outcomes. The applications include, but are not limited to, cancer detection [80, 81], skin segmentation [82], and pill identification [83, 84] which showcase the model’s ability to adapt to different needs, and essential tasks.
In industrial applications, YOLO aids in surface inspection processes to detect defects and anomalies [64, 65], structural health monitoring[85], reliability and safety[86] ensuring quality control in manufacturing and production[64].
Surveillance and security systems also leverage YOLO for real-time monitoring and rapid identification of suspicious activities [66, 67]. By integrating these models into surveillance systems, security personnel can more effectively monitor and respond to potential threats, enhancing public safety [87]. Similarly, in the context of public health measures like social distancing and face mask detection during pandemics [88, 89], YOLO models provided essential support in enforcing health regulations.
In agriculture, YOLO models have been applied to detect and classify crops [90], pests, and diseases [91, 92, 93], facilitating precision agriculture techniques and automating farming operations to increase productivity and optimizing inputs. Additionally, in remote sensing, YOLO contributes to object recognition in satellite [94, 95] and aerial imagery [96, 97], which supports urban planning, land use mapping, and environmental monitoring. These capabilities demonstrate YOLO’s contribution to critical global challenges such as urban development and environmental conservation.
@@ III Review of YOLO Versions
This section reviews YOLO series models, starting from the advanced and latest version, YOLO11, and progressively tracing back to the foundational YOLOv1. By first highlighting the most recent technological advancements, this approach enables immediate insights into the state-of-the-art capabilities of object detection. Subsequently, the narrative explores how earlier models laid the groundwork for these innovations.
III-A YOLOv11 and it’s Predecessors
@@
YOLOv11, developed by Ultralytics represents the most recent version building upon the foundations established by its predecessors in the YOLO family. This latest iteration introduces several architectural innovations that enhance its performance across a wide spectrum of tasks as depicted in Figure 4. The model incorporates the C3k2 (Cross Stage Partial with kernel size 2) block, which replaces the C2f block used in previous versions, offering improved computational efficiency [128]. Additionally, YOLOv11 retains the SPPF (Spatial Pyramid Pooling - Fast) component and introduces the C2PSA (Convolutional block with Parallel Spatial Attention) module, collectively enhancing feature extraction capabilities [78]. These architectural enhancements enable YOLOv11 to capture intricate image details with greater precision, particularly in challenging scenarios involving small or occluded objects. The model’s versatility is evident in its support for a broad range of computer vision tasks, including object detection, instance segmentation, pose estimation, image classification, and oriented bounding box (OBB) detection (https://docs.ultralytics.com/models/yolo11/).
@@
Empirical evaluations of YOLOv11 demonstrate its superior performance in terms of both accuracy and efficiency. Notably, the YOLOv11m variant achieves a higher mean Average Precision (mAP) score on the COCO dataset while utilizing 22% fewer parameters compared to its YOLOv8m counterpart127. This reduction in parameter count contributes to faster model performance without significantly impacting overall accuracy. Furthermore, YOLOv11 exhibits inference times approximately 2% faster than YOLOv10, making it particularly well-suited for real-time applications37. The model’s efficiency extends across various deployment environments, including edge devices, cloud platforms, and systems supporting NVIDIA GPUs. YOLOv11 is available in multiple variants, ranging from nano to extra-large, catering to diverse computational requirements and use cases. These advancements position YOLOv11 as a state-of-the-art solution for industries requiring rapid and accurate image analysis, such as autonomous driving, surveillance, and industrial automation.
@@
YOLOv10, the penultimate in the series, incorporates advanced techniques like automated architecture search and more refined loss functions to enhance detection accuracy and speed, tailored for both edge and cloud computing environments. This version and its predecessors, YOLOv9 and YOLOv8, introduce substantial improvements in network architecture, such as the integration of cross-stage partial networks (CSPNets) and the use of transformer-based backbones for better feature extraction across different scales. YOLOv7 and YOLOv6 continued to build on these improvements by optimizing computational efficiency and expanding model scalability. Meanwhile, YOLOv5 introduced PyTorch support, which significantly enhanced the model’s accessibility and adaptability, thus broadening its application in industry and academia. YOLOv4 marked a pivotal point by integrating features like Mish activation and Cross-Stage Partial connections, setting new standards for speed and accuracy in real-time applications. The mid-generations, starting with YOLOv3, were notable for introducing multi-scale predictions and bounding box predictions across different layers, which greatly improved the model’s ability to detect small objects—a longstanding challenge in earlier versions. YOLOv3 was also one of the first to leverage deeper feature extractors like Darknet-53, which significantly boosted its performance over YOLOv2. YOLOv2 itself had introduced important features such as batch normalization and high-resolution classifiers, which enhanced the overall accuracy without compromising the speed. The original YOLO model, YOLOv1, was revolutionary, proposing a single-stage detection framework that unified the detection process into a single neural network model.
Figure 5 a illustrates a sophisticated transformer-based model that simplifies the detection process by integrating dual label assignments and eliminating the need for non-max suppression (NMS), achieving a streamlined, end-to-end object detection. YOLOv9, as shown in Figure 5 b, introduces the Programmable Gradient Information (PGI) system to enhance model interpretability and robustness, significantly improving generalization across various tasks. Moving to YOLOv8 and YOLOv7, Figures 5 c and 5 d respectively depict their architectures which incorporate elements like ELAN and CSPNet to boost performance and flexibility across computing devices. YOLOv6, highlighted in Figure 5 e, focuses on industry applications with enhancements in model quantization and real-time performance.
@@ III-B YOLOv10, YOLOv9 and YOLOv8
YOLOv10 [112], developed at Tsinghua University, China, represents a breakthrough in the YOLO series for real-time object detection, achieving unprecedented performance. This version eliminates the need for non-maximum suppression (NMS)[131], a traditional bottleneck in earlier models, thereby drastically reducing latency. YOLOv10 introduces a dual assignment strategy in its training protocol, which optimizes detection accuracy without sacrificing speed with the help of one-to-many and one-to-one label assignments, ensuring robust detection with lower latency[132, 133]. The architecture of YOLOv10 includes several innovative components that enhance both computational efficiency and detection performance. Among these are lightweight classification heads[134] that reduce computational demands, spatial-channel decoupled downsampling to minimize information loss during feature reduction[135], and rank-guided block design that optimizes parameter use[136]. These architectural advancements ensure that YOLOv10 operates synergistically across various scales—from YOLOv10-N (Nano) to YOLOv10-X (Extra Large), making it adaptable to diverse computational constraints and operational requirements [112]. According to wang et al. [112], performance evaluations on benchmark datasets like MS-COCO[137] demonstrate that YOLOv10 not only surpasses its predecessors—YOLOv9 and YOLOv8—in both accuracy and efficiency but also sets new industry standards. For instance, YOLOv10-S substantially outperforms comparable models (e.g., YOLOv9 BASE, yoloV9 Gelan, YOLOv8, YOLOv7) with an improved mAP and lower latency. This version also incorporates holistic efficiency-accuracy driven design, large-kernel convolutions, and partial self-attention modules, collectively improving the trade-off between computational cost and detection capability. The architecture diagrams of YOLOv10, YOLOv9, and YOLOv8 are summarized in Figures 6, 7, and 8, respectively.

### v6

# YOLOv12 to Its Genesis: A Decadal and Comprehensive Review of The You Only Look Once (YOLO) Series

@@ Abstract
This review systematically examines the progression of the You Only Look Once (YOLO) object detection algorithms from YOLOv1 to the recently unveiled YOLOv12. Employing a reverse chronological analysis, this study examines the advancements introduced by YOLO algorithms, beginning with YOLOv12 and progressing through YOLO11 (or YOLOv11), YOLOv10, YOLOv9, YOLOv8, and subsequent versions to explore each version’s contributions to enhancing speed, detection accuracy, and computational efficiency in real-time object detection. Additionally, this study reviews the alternative versions derived from YOLO architectural advancements of YOLO-NAS, YOLO-X, YOLO-R, DAMO-YOLO, and Gold-YOLO. By detailing the incremental technological advancements in subsequent YOLO versions, this review chronicles the evolution of YOLO, and discusses the challenges and limitations in each of the earlier versions. The evolution signifies a path towards integrating YOLO with multimodal, context-aware, and Artificial General Intelligence (AGI) systems for the next YOLO decade, promising significant implications for future developments in AI-driven applications.
@@
You Only Look Once, YOLO, YOLOv1 to YOLOv12, YOLOv11, YOLOv10, YOLO configurations, CNN, Deep learning, Real-time object detection, Artificial intelligence, Computer vision, Healthcare and Medical Imaging, Autonomous vehicles, Traffic safety, Industrial manufacturing, Surveillance, Agriculture
@@ I-D You Only Look Once Approach
The ”You Only Look Once” (YOLO) object detection algorithm was first introduced by Joseph Redmon et al., [58] in 2015, revolutionized real-time object detection by combining region proposal and classification into a single neural network, significantly reducing computation time. YOLO’s unified architecture divides the image into a grid, predicting bounding boxes and class probabilities directly for each cell, enabling end-to-end learning [58]. YOLOv1 utilized a simplified CNN backbone, setting the stage with basic bounding box predictions. Subsequent versions like YOLOv2 incorporated the DarkNet-19 backbone and refined anchor boxes through K-means clustering. YOLOv3 expanded this further with a DarkNet-53 architecture, integrating multi-scale detection and residual connections. The series continued to innovate with YOLOv4 implementing CSPDarkNet-53 and PANet alongside mosaic data augmentation. YOLOv5 and YOLOv6 introduced CSPNet with dynamic anchor refinement and further enhancements in PANet, respectively. YOLOv7 featured an EfficientRep backbone with dynamic label assignment, while YOLOv8 introduced a Path Aggregation Network with Dynamic Kernel Attention. YOLOv9 developed multi-level auxiliary feature extraction, and YOLOv10 optimized the system with a lightweight classification head and distinct spatial and channel transformations. YOLOv11 introduced the C3k2 block in its backbone and utilized C2PSA for improved spatial attention. The latest, YOLOv12, marks a significant shift towards an attention-centric design, introducing the Area Attention (A2) module for efficient large receptive field processing, Residual Efficient Layer Aggregation Networks (R-ELAN) for enhanced feature aggregation, and architectural optimizations including FlashAttention and adjusted MLP ratios. This attention-based approach allows YOLOv12 to achieve state-of-the-art performance in both accuracy and efficiency, surpassing previous CNN-based models while maintaining real-time detection capabilities [59]. autonomous vehicles and traffic safety [4, 60, 16, 61], healthcare [62, 63, 64], industrial applications [65, 66, 60], surveillance and security [67, 68] and agriculture [69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79], where accuracy and speed are crucial.
@@ III Review of YOLO Versions
This section reviews YOLO series models, starting from the advanced and latest version, YOLOv12, and progressively tracing back to the foundational YOLOv1. By first highlighting the most recent technological advancements, this approach enables immediate insights into the state-of-the-art capabilities of object detection. Subsequently, the narrative explores how earlier models laid the groundwork for these innovations.
III-A YOLOv12 and YOLO11
YOLOv12 [59] is the most recent YOLO version introduced in February 2025, which marks a substantial advancement in real-time object detection by integrating attention mechanisms into the YOLO framework while maintaining competitive inference speeds. This attention-centric framework not only surpasses popular real-time detectors (e.g., xxx) in accuracy but also achieves state-of-the-art performance through a combination of innovative attention methods, Residual Efficient Layer Aggregation Networks (R-ELAN), and several architectural optimizations.
Figure 4: (a) Latency comparison on the MS COCO benchmark reveals significantly faster inference achieved with YOLOv12 compared to the same achieved with previous YOLO versions. (b) GFLOPs analysis also shows enhanced computational efficiency. [59]
Figure 4a demonstrates latency comparisons on the MS COCO benchmark dataset, highlighting YOLOv12’s significantly lower inference latency compared to YOLOv11, YOLOv10, YOLOv9, and YOLOv8. The curve reveals a substantial reduction in latency, enabling faster processing speeds while maintaining high detection accuracy (reference). Complementing this, Figure 4b presents comparison in terms of GFLOPs, which shows YOLOv12 achieves higher computational efficiency, reflecting its ability to handle complex computations effectively. This balance between speed and computational power demonstrates YOLOv12’s robust performance.
On the COCO dataset, YOLOv12 sets a new state-of-the-art standard, with the lightweight YOLOv12-N achieving 40.6% mAP and the larger YOLOv12-X reaching 55.2% mAP. These results, combined with the latency and GFLOPs improvements, establish YOLOv12 as a new benchmark in real-time object detection. The model framework is available in five scales: YOLOv12-N, S, M, L, and X, each optimized for specific applications. For instance, YOLOv12-N achieves 40.6% mAP at 1.64 ms on a T4 GPU, outperforming YOLOv10-N and YOLOv11-N by 2.1% and 1.2% mAP, respectively. Similarly, YOLOv12-S attains 48.0% mAP at 2.61 ms/image, surpassing YOLOv8-S, YOLOv9-S, YOLOv10-S, and YOLOv11-S by margins of 3.0%, 1.2%, 1.7%, and 1.1% mAP. The larger models in the YOLOv12 family continue to show improvements in efficiency and performance. Notably, YOLOv12-M achieves 52.5% mAP at 4.86 ms/image. In terms of computational effieciency, YOLOv12-L demonstrates a significant reduction in FLOPs, decreasing by 31.4G compared to its predecessor, YOLOv10-L. At the highest end of the scale, YOLOv12-X showcases superior performance, outperforming both YOLOv10-X and YOLOv11-X in detection accuracy [59].
YOLOv12 also surpasses end-to-end detectors like RT-DETR and RT-DETRv2. For example, YOLOv12-S runs 42% faster than RT-DETR-R18 and RT-DETRv2-R18, using only 36% of the computation and 45% of the parameters. Residual connections show minimal impact on convergence in smaller models (YOLOv12-N) but are critical for stable training in larger models (YOLOv12-L/X), with YOLOv12-X requiring a scaling factor of 0.01. The area attention module reduces inference time by 0.7 ms on an RTX 3080 with FP32 precision, while FlashAttention further accelerates inference by 0.3–0.4 ms.
Visualization analyses confirm that YOLOv12 produces clearer object contours and more precise foreground activations than its predecessors. A convolution-based attention implementation proves to be faster than linear alternatives. Additionally, a hierarchical design, extended training (approximately 600 epochs), an optimized convolution kernel size ($7\times 7$), absence of positional embedding, and an MLP ratio of 1.2 collectively enhance the framework’s performance and efficiency.
Figure 5: (a) Architecture of the YOLOv12 object detection model that integrated Area Attention (A²) modules, R-ELAN blocks, and a streamlined detection head; (b) Comparison of ”Attention Module” architectures: CSPNet [110], ELAN [111], C3K2 (used in YOLOv9) [91], and the novel R-ELAN [59] introduced with YOLOv12, which improved residual connections and enhanced feature aggregation, demonstrating superior performance.
As discussed before, the YOLOv12 architecture (Figure 5a), demonstrates an advanced integration of A2 (Area Attention) modules, R-ELAN (Residual Efficient Layer Aggregation Networks) blocks, and a streamlined detection head. This design optimizes the model’s visual information processing while maintaining high accuracy. The major innovations on the YOLOv12 architecture are listed below.
III-A1 YOLOv12 Architectural Innovation
  • Area Attention (A2) Module: This module implements segmented feature processing with Flash Attention integration, reducing computational complexity by 50% through spatial reshaping while maintaining large receptive fields. AA enables real-time detection at fixed $n=640$ resolution through optimized memory access patterns, as illustrated in Figure 5a.
  • Residual ELAN (R-ELAN) Hierarchy: R-ELAN combines residual shortcuts (0.01 scaling) with dual-branch processing to mitigate the gradient vanishing problem. The model also features a streamlined final aggregation stage that reduces parameters by 18% and FLOPs by 24% compared to baseline architectures, as shown in Figure 5b.
  • Efficient Architectural Revisions: YOLOv12 replaces positional encoding with 7×7 depth-wise convolution for implicit spatial awareness. It also implements adaptive MLP ratio (1.2×) and shallow block stacking to balance the computational load, achieving 4.1 ms inference latency on V100 hardware.
  • Optimized Training Framework: The model was trained over 600 epochs using SGD with cosine scheduling (initial lr=0.01). The model also incorporates Mosaic-9 and Mixup augmentations with 12.8% mAP gain on COCO dataset, maintaining real-time performance through selective kernel convolution integration.
Figure 5b presents an architectural comparison of popular attention modules: CSPNet, ELAN, C3K2 (a case of GELAN), and the proposed R-ELAN. Brief summary of these modules is blow.
  • CSPNet (Cross Stage Partial Network): CSPNet enhances gradient flow by splitting feature maps into two paths, one for learning and one for propagation, reducing computational bottlenecks and improving inference speed. This model is visually depicted in Figure 5b (leftmost module).
  • ELAN (Efficient Layer Aggregation Network): ELAN improves feature integration by aggregating multi-scale features efficiently, enhancing the model’s ability to detect objects at various scales. However, as shown in Figure 5b (second module), ELAN can introduce instability due to gradient blocking and lacks of residual connections, particularly in large-scale models.
  • C3K2 (Compact GELAN): This module is a compact version of GELAN (Generalized Efficient Layer Aggregation Network) that offers a balance between computational efficiency and feature expressiveness, suitable for resource-constrained environments. The module is also illustrated in Figure 5b (third module).
  • R-ELAN (Residual ELAN): R-ELAN introduces residual connections and redesigns feature aggregation to address optimization challenges in attention-based models, combining the benefits of residual learning with efficient feature aggregation. As shown in Figure 5b (rightmost module), R-ELAN applies a residual shortcut with a scaling factor (default 0.01) and processes the input through a transition layer, followed by a bottleneck structure for improved stability and performance.
The R-ELAN design, as depicted in Figure 5b, addresses the limitations of ELAN by introducing residual connections and a revised aggregation approach. Unlike ELAN, which splits the input into two parts and processes them separately, R-ELAN applies a transition layer to adjust channel dimensions and processes the feature map through subsequent blocks before concatenation. This design mitigates gradient blocking and ensures stable convergence, particularly in large-scale models like YOLOv12-L and YOLOv12-X. The integration of residual connections and attention mechanisms in R-ELAN, as shown in Figure 5b, highlights YOLOv12’s architectural advancements in balancing efficiency and accuracy.
@@ III-A1 YOLOv12 Architectural Innovation
YOLO11: YOLO11 developed by Ultralytics represents the most recent version building upon the foundations established by its predecessors in the YOLO family. This latest iteration introduces several architectural innovations that enhance its performance across a wide spectrum of tasks as depicted in Figure 6. The model incorporates the C3k2 (Cross Stage Partial with kernel size 2) block, which replaces the C2f block used in previous versions, offering improved computational efficiency [112]. Additionally, YOLOv11 retains the SPPF (Spatial Pyramid Pooling - Fast) component and introduces the C2PSA (Convolutional block with Parallel Spatial Attention) module, collectively enhancing feature extraction capabilities [79]. These architectural enhancements enable YOLOv11 to capture intricate image details with greater precision, particularly in challenging scenarios involving small or occluded objects. The model’s versatility is evident in its support for a broad range of computer vision tasks, including object detection, instance segmentation, pose estimation, image classification, and oriented bounding box (OBB) detection (https://docs.ultralytics.com/models/yolo11/).
@@
Empirical evaluations of YOLOv achieves a higher mean Average Precision (mAP) score on the COCO dataset while utilizing 22% fewer parameters compared to its YOLOv8m counterpart [113]. This reduction in parameter count contributes to faster model performance without significantly impacting overall accuracy. Furthermore, YOLOv11 exhibits inference times approximately 2% faster than YOLOv10, making it particularly well-suited for real-time applications. The model’s efficiency extends across various deployment environments, including edge devices, cloud platforms, and systems supporting NVIDIA GPUs. YOLOv11 is available in multiple variants, ranging from nano to extra-large, catering to diverse computational requirements and use cases. These advancements position YOLOv11 as a state-of-the-art solution for industries requiring rapid and accurate image analysis, such as autonomous driving, surveillance, and industrial automation.
@@
YOLOv10, incorporates advanced techniques like automated architecture search and more refined loss functions to enhance detection accuracy and speed, tailored for both edge and cloud computing environments. This version and its predecessors, YOLOv9 and YOLOv8, introduce substantial improvements in network architecture, such as the integration of cross-stage partial networks (CSPNets) and the use of transformer-based backbones for better feature extraction across different scales. YOLOv7 and YOLOv6 continued to build on these improvements by optimizing computational efficiency and expanding model scalability. Meanwhile, YOLOv5 introduced PyTorch support, which significantly enhanced the model’s accessibility and adaptability, thus broadening its application in industry and academia. YOLOv4, on the other hand, marked a pivotal point in YOLO history by integrating features like Mish activation and Cross-Stage Partial connections, setting new standards for speed and accuracy in real-time applications. The mid-generations, starting with YOLOv3, were notable for introducing multi-scale predictions and bounding box predictions across different layers, which greatly improved the model’s ability to detect small objects—a longstanding challenge in earlier versions. YOLOv3 was also one of the first YOLO models to leverage deeper feature extractors like Darknet-53, which significantly boosted its performance over YOLOv2. YOLOv2 itself had introduced important features such as batch normalization and high-resolution classifiers, which enhanced the overall accuracy without compromising the speed. The original YOLO model, YOLOv1, was revolutionary, proposing a single-stage detection framework that unified the object detection process into a single neural network model.
Figure 7 a illustrates a sophisticated transformer-based model that simplifies the detection process by integrating dual label assignments and eliminating the need for non-max suppression (NMS), achieving a streamlined, end-to-end object detection. YOLOv9, as shown in Figure 7 b, introduces the Programmable Gradient Information (PGI) system to enhance model interpretability and robustness, significantly improving generalization across various tasks. Moving to YOLOv8 and YOLOv7, Figures 7 c and 7 d respectively depict their architectures which incorporate elements like ELAN and CSPNet to boost performance and flexibility across computing devices. YOLOv6, highlighted in Figure 7 e, focuses on industry applications with enhancements in model quantization and real-time performance.
@@ III-B YOLOv10, YOLOv9 and YOLOv8
YOLOv10 [94], developed at Tsinghua University, China, represents a breakthrough in the YOLO series for real-time object detection, achieving unprecedented performance. This version eliminates the need for non-maximum suppression (NMS)[116], a traditional bottleneck in earlier models, thereby drastically reducing latency. YOLOv10 introduces a dual assignment strategy in its training protocol, which optimizes detection accuracy without sacrificing speed with the help of one-to-many and one-to-one label assignments, ensuring robust detection with lower latency[117, 118]. The architecture of YOLOv10 includes several innovative components that enhance both computational efficiency and detection performance. Among these are lightweight classification heads[119] that reduce computational demands, spatial-channel decoupled downsampling to minimize information loss during feature reduction[120], and rank-guided block design that optimizes parameter use[121]. These architectural advancements ensure that YOLOv10 operates synergistically across various scales—from YOLOv10-N (Nano) to YOLOv10-X (Extra Large), making it adaptable to diverse computational constraints and operational requirements [94]. According to wang et al. [94], performance evaluations on benchmark datasets like MS-COCO[122] demonstrate that YOLOv10 not only surpasses its predecessors—YOLOv9 and YOLOv8—in both accuracy and efficiency but also sets new industry standards. For instance, YOLOv10-S substantially outperforms comparable models (e.g., YOLOv9 BASE, yoloV9 Gelan, YOLOv8, YOLOv7) with an improved mAP and lower latency. This version also incorporates holistic efficiency-accuracy driven design, large-kernel convolutions, and partial self-attention modules, collectively improving the trade-off between computational cost and detection capability. The architecture diagrams of YOLOv10, YOLOv9, and YOLOv8 are summarized in Figures 8, 9, and 10, respectively.